# SurvReLU: Inherently Interpretable Survival Analysis via Deep ReLU Networks

### v1

@@ Abstract.
Survival analysis models time-to-event distributions with censorship. Recently, deep survival models using neural networks have dominated due to their representational power and state-of-the-art performance. However, their ”black-box” nature hinders interpretability, which is crucial in real-world applications. In contrast, ”white-box” tree-based survival models offer better interpretability but struggle to converge to global optima due to greedy expansion. In this paper, we bridge the gap between previous deep survival models and traditional tree-based survival models through deep rectified linear unit (ReLU) networks. We show that a deliberately constructed deep ReLU network (SurvReLU) can harness the interpretability of tree-based structures with the representational power of deep survival models. Empirical studies on both simulated and real survival benchmark datasets show the effectiveness of the proposed SurvReLU in terms of performance and interoperability. The code is available at https://github.com/xs018/SurvReLU.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/XXXXXX.XXXXXXccs: Computing methodologies Machine learningccs: Computing methodologies Machine learning algorithmsccs: Computing methodologies Neural networks
@@ 1. Introduction
Tree-based survival models (Gordon and Olshen, 1985; Kogalur et al., 2008) are competing approaches to deep survival models. These models either minimize local homogeneity (Gordon and Olshen, 1985) (measured by log-rank test) or maximize local heterogeneity (Wang et al., 2016) (measured by Kullback–Leibler divergence) between branches at each splitting node. (Kretowska, 2019) then proposes an oblique survival tree using a non-axis-aligned splitting rule as opposed to axis-aligned splitting rules in (Gordon and Olshen, 1985; Kogalur et al., 2008)(Kogalur et al., 2008; Jaeger et al., 2019) extend these concepts to random (oblique) survival forests to enhance their generalizability. Although these ”white-box” tree-based survival models offer more interpretability compared to deep survival models, they are typically shown to be inferior to deep survival models (Katzman et al., 2018; Kvamme et al., 2019; Lee et al., 2018; Ren et al., 2019). This may be due to their inability to guarantee convergence to global optima because of their greedy expansion and reliance on predefined splitting rules.
@@
In this paper, we show that a thoughtfully designed ReLU network can achieve interpretability, by forming a tree structure, and maintain the representational power of a neural network. Specifically, we establish an explicit connection between the ReLU network and previous tree-based survival models, leveraging ReLU networks’ ability to partition the input space into locally homogeneous regions like a tree (Sec. 2.1). Second, we introduce a novel statistically-driven method to dynamically optimize the topology of the ReLU network specifically for survival analysis (Sec. 2.2). This can help compact the tree structure and improve the interpretability. To this end, we construct the novel ReLU network that is inherently interpretable like a tree (termed SurvReLU). Third, we show that SurvReLU can be optimized end-to-end with a flexible choice of loss functions (Sec. 2.3). Experiments on both simulated and real datasets demonstrate the effectiveness of the proposed SurvReLU in terms of both performance and interpretability.
@@ 2. Method
The right-censored survival data is given in the form of $N$ triples $\{\boldsymbol{x},T,E\}_{n=1}^{N}$, where $\boldsymbol{x},T,E$ denote the $d-$dimensional covariates, the time of record, and an event indicator, respectively. With the random variable $t$ representing time, the objective is to model the probability of the event occurring over time $t$. This is commonly formulated to compute the survival function $S(t)$: the probability that an event occurs beyond time $t$,
@@ 2.1. Network Architecture
where $\texttt{concat}[\cdot]$ denotes concatenation operation. The resulting network architecture is shown in Fig. 1(a). This modification enables the partitioning of polyhedrons at each layer to resemble a non-axis-aligned tree-splitting rule, where the absolute value of the weight in $\boldsymbol{W}^{(l)}$ indicates the importance of each covariate at current partitioning. It is worth noting that a sparse weight matrix $\boldsymbol{W}^{(l)}$ (i.e., $\boldsymbol{W}^{(l)}$ has only one nonzero entry) can also lead to an axis-aligned splitting rule. Up to now, we have established the explicit connection between the ReLU network and previous axis-aligned (Gordon and Olshen, 1985; Wang et al., 2016) and non-axis-aligned (Jaeger et al., 2019) survival trees. Since the derived ReLU network itself is a neural network, its connection to other networks (e.g., standard MLPs) used in (Katzman et al., 2018; Kvamme et al., 2019; Lee et al., 2018; Ren et al., 2019) is trivial.
Figure 2. The change of rank ratio of matrix $\boldsymbol{O}$ over time.
@@ 2.2. Statistically-Driven Topology Optimization
Although we have constructed a ReLU network whose behavior inherently resembles a tree, it is not specifically optimized for survival analysis. One notable limitation of the previously derived ReLU network in Sec. 2.1 is that its network topology is defined as a priori. This implies that the tree constructed by the ReLU network is fully expanded (i.e., a dense tree).
As the activation patterns $\boldsymbol{O}$ uniquely encoded the partitioning of input space in the derived ReLU network, the topology of the learned tree structure can be reflected by $\boldsymbol{O}$. We observe that the activation patterns matrix $\boldsymbol{O}$ is typically low-ranked (see Fig. 2). This suggests that the learned ReLU network, however, produces a sparse tree topology: some partitions/nodes should be pruned. Based on this observation, we employ the log-rank test to localize which partition should be pruned. Specifically, at each forward pass, we can reconstruct the tree structure of the derived ReLU network recursively (see Fig. 1(c)). Simultaneously, we can perform the log-rank test at each splitting node and merge the nodes whose p-value from the log-rank test is less than 0.05 ($p<0.05$) by setting the corresponding activation patterns to 0. This enables us to dynamically optimize the derived ReLU network’s topology and automatically prune the resulting tree structures.
But when to perform the proposed topology optimization remains an open challenge. Model performance may deteriorate when optimizing network topology after training (Molchanov et al., 2016). However, optimizing network topology at early training iterations may restrict its representational capacity (Blalock et al., 2020). Again, we can leverage the rank of the activation patterns $\boldsymbol{O}$ to decide when to optimize the network topology. To achieve this, we keep track of the changes in the rank of the matrix $\boldsymbol{O}$ over time. Once the topology of the learned tree structures is unchanged, we optimize the network topology.
@@ 2.3. Model Learning
Complexity analysis. Compared to standard MLP, SurvReLU computes the derivative of each neuron w.r.t. $\boldsymbol{x}$, which requires an additional back-propagation (i.e., $\mathcal{O}(L^{3})$). By utilizing the dynamic programming trick outlined in (Lee and Jaakkola, 2020), this can be reduced to $\mathcal{O}(L^{2})$. The log-rank test also introduces additional computation, which will take $\mathcal{O}(N\log N)$ to sort the time $T$. However, this can be handled once at the very beginning of each training iteration.
@@
4:     reconstruct the tree node at $l$-th layer (see Fig. 1(c))
5:     log-rank test on two partitions ($\boldsymbol{z^{(l)}}\geq 0$ and $\boldsymbol{z^{(l)}}<0$)
@@
Methods Simulated Linear Simulated Gaussian SUPPORT METABRIC
@@
3. Experiments and Results
We conducted experiments on two simulated and real datasets.
Simulated:
@@ 3. Experiments and Results
The simulated datasets were sourced from (Katzman et al., 2018), consisting of a training ($N=4000$), validation ($N=1000$), and testing ($N=1000$) set. Each sample includes $d=10$ covariates (i.e., $x_{0},x_{1},\cdots,x_{9}$) drawn from a uniform distribution on $[-1,1)$. The event time $T$ was generated based on an exponential Cox model. Further information about the simulation process can be found in (Katzman et al., 2018). These simulated datasets encompass two subcategories characterized by either a linear risk function (i.e., $x_{1}$: $h(x)=x_{0}+2x_{1}$) or a Gaussian risk function (i.e., $h(\boldsymbol{x})=\log(\lambda_{max})\exp(-(x_{0}^{2}+x_{1}^{2})/2r^{2})$; $\lambda_{max}=5.0$ and $r=0.5$).
Real:
Two popular benchmark datasets were used: (i) SUPPORT dataset, originating from the Study to Understand Prognoses Preference Outcomes and Risks of Treatment (Knaus et al., 1995). It comprises 8873 patients (patients with missing features were dropped) with 14 covariates. The censorship rate within this dataset stands at approximately $32\%$. (ii) METABRIC dataset is designed to improve breast cancer treatment recommendations (Curtis et al., 2012), including gene expression and clinical data for 1,980 patients. The data was processed to integrate four gene indicators and five clinical features, resulting in a total of 9 covariates. The proposition of censorship is around $42\%$.
3.1. Experiments
Evaluation metrics.
@@ 3.1. Experiments
The evaluation of the performance for survival analysis was carried out by the time-dependent concordance index ($C^{td}_{\tau}$). Specifically, the C-index is designed to assess a model’s ability to predict the ordering of event times through ranking.
Experimental setup.
@@
(i) On the simulated datasets, we adhere to the default training, validation, and testing data splits outlined in (Katzman et al., 2018). All the covariates are z-scored to have a zero mean and unit variance. The 95% confidence intervals are obtained by bootstrapping (Katzman et al., 2018). (ii) We conduct five-fold cross-validation for the real datasets (i.e., SUPPORT and METABRIC) to thoroughly assess the performance and generalizability of all methods. The mean ($\pm$ standard deviation) of C-index is reported. For preprocessing, we apply one-hot encoding to all categorical variables and z-score for all numerical variables. (iii) The hyperparameter tuning for the proposed model (i.e., learning rate, batch size, and $L$) is carried out by the random search (Bergstra and Bengio, 2012). The optimal learning rate is found to be $0.1$. The batch size is set to 1024. We recommend a large batch size to ensure stable statistics in the log-rank test. The optimal number of layers is [6, 30, 16, 14] for simulated Linear, simulated Gaussian, SUPPORT, and METABRIC, respectively. All the experiments are implemented in PyTorch and performed on a Nvidia GTX 1080Ti.
3.2. Results
@@
Main results. As shown in Table 1, the proposed SurvReLU competes favorably with other deep survival models and tree-based survival models on both simulated and real datasets in terms of $C^{td}$. We observe that (i) on the simulated datasets, SurvReLU(Cont.) and SurvReLU(Cont.) achieve similar performance but still outperform other competing methods. Remarkably, SurvReLU can approximate the true Linear/Gaussian risk function as effectively as other deep survival models (see Fig. 3(a)). However, the other axis-aligned and even non-axis-aligned tree-based survival models show inferior performance in modeling the risk function. This suggests that SurvReLU retains the representational capacity of a neural network. Although SurvReLU has a non-axis-aligned partitioning rule similar to ORSF, we conjecture that the superiority of SurvReLU over other tree-based survival models arises from its end-to-end optimization instead of greedy expansion (e.g., fitting a CPH at each split in ORSF). (ii) On the real datasets, SurvReLU (Cont.) outperforms other competing continuous-time deep survival models (DeepSurv and Cox-CC). SurvReLU (Disc.) achieves even better performance than discrete-time DeepHit and outperforms other tree-based survival models, even without the use of ensembles. Importantly, SurvReLU results in a compact tree structure that is inherently as interpretable as tree-based survival models (Fig. 3(b) and (c)), a feature none of the other deep survival models can provide.
@@
Ablation studies. We conduct ablation on the number of layers $L$ in a SurvReLU network as well the sparsity of the weight matrices $\boldsymbol{W}=\{\boldsymbol{W}^{(1)},\cdots,\boldsymbol{W}^{(L)}\}$. The sparsity of $\boldsymbol{W}$ is enforced by recursively soft-thresholding $\boldsymbol{W}$ with a strength $\lambda$ (Kusupati et al., 2020; Donoho, 1995; Qiu et al., 2023). We observed that (i) similar to many deep survival models, the performance of SurvReLU typically saturates after reaching a certain number of layers (e.g., $L=20$ in Fig. 4(a)). (ii) The performance of SurvReLU is traded-off with the sparsity of $\boldsymbol{W}$ and hence the interpretability of the model (Fig. 4(b)). This is because a more sparse $\boldsymbol{W}$ pushes SurvReLU to result in a tree structure with an axis-aligned partitioning rule (see Sec. 2.1), which is better to interpret.
@@
Figure 4. Ablations on (a) number of layers $L$ and (b) sparsity of weight matrix $\boldsymbol{W}$ using the simulated Gaussian dataset.

### v2

@@ Abstract.
Survival analysis models time-to-event distributions with censorship. Recently, deep survival models using neural networks have dominated due to their representational power and state-of-the-art performance. However, their ”black-box” nature hinders interpretability, which is crucial in real-world applications. In contrast, ”white-box” tree-based survival models offer better interpretability but struggle to converge to global optima due to greedy expansion. In this paper, we bridge the gap between previous deep survival models and traditional tree-based survival models through deep rectified linear unit (ReLU) networks. We show that a deliberately constructed deep ReLU network (termed SurvReLU) can harness the interpretability of tree-based structures with the representational power of deep survival models. Empirical studies on both simulated and real survival benchmark datasets showed the effectiveness of the proposed SurvReLU in terms of performance and interoperability. The code is available at  https://github.com/xs018/SurvReLU.
@@
copyright: acmlicensedjournalyear: 2024isbn: 978-1-4503-XXXX-X/18/06journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679947ccs: Computing methodologies Machine learningccs: Computing methodologies Machine learning algorithmsccs: Computing methodologies Neural networks
@@ 1. Introduction
Tree-based survival models (Gordon and Olshen, 1985; Kogalur et al., 2008) are competing approaches to deep survival models. These models either minimize local homogeneity (Gordon and Olshen, 1985) (measured by log-rank test) or maximize local heterogeneity (Wang et al., 2016) (measured by Kullback–Leibler divergence) between branches at each splitting node. Oblique survival trees (Kretowska, 2019; Bertsimas et al., 2022; Huisman et al., 2024; Zhang et al., 2024), which use non-axis-aligned splitting rules as opposed to axis-aligned survival trees (Gordon and Olshen, 1985), have also been explored. (Kogalur et al., 2008; Jaeger et al., 2019) extend the concepts of survival trees to random (oblique) survival forests to enhance their generalizability. Although these ”white-box” tree-based survival models offer more interpretability compared to deep survival models, they are typically shown to be inferior to deep survival models (Katzman et al., 2018; Kvamme et al., 2019; Lee et al., 2018; Ren et al., 2019). This may be due to their inability to guarantee convergence to global optima because of their greedy expansion and reliance on predefined splitting rules.
@@
In this paper, we show that a thoughtfully designed ReLU network can achieve interpretability, by forming a tree structure, and maintain the representational power of a neural network. Specifically, we establish an explicit connection between the ReLU network and previous tree-based survival models, leveraging ReLU networks’ ability to partition the input space into locally homogeneous regions like a tree (Sec. 2.1). Second, we introduce a novel statistically-driven method to dynamically optimize the topology of the ReLU network specifically for survival analysis (Sec. 2.2). This can help compact the tree structure and improve the interpretability. To this end, we construct the novel ReLU network that is inherently interpretable like a tree (termed SurvReLU). Third, we show that SurvReLU can be optimized end-to-end with a flexible choice of loss functions (Sec. 2.3). Experiments on both simulated and real datasets demonstrated the effectiveness of the proposed SurvReLU in terms of both performance and interpretability.
@@ 2. Method
The right-censored survival data is given in the form of $N$ triples $\{\boldsymbol{x},T,E\}_{n=1}^{N}$, where $\boldsymbol{x},T,E$ denote the $d-$dimensional covariates, the time of record, and an event indicator, respectively. With the random variable $t$ representing time, the objective is to model the probability of the event occurring over time $t$. This is commonly formulated to compute the survival function $S(t)$, i.e., the probability that an event occurs beyond time $t$:
@@ 2.1. Network Architecture
where $\texttt{concat}[\cdot]$ denotes concatenation operation. The resulting network architecture is shown in Fig. 1(a). This modification enables the partitioning of polyhedrons at each layer to resemble a non-axis-aligned tree-splitting rule, where the absolute value of each entry in $\boldsymbol{W}^{(l)}$ indicates the importance of each covariate at the $l$-th layer. Note that a sparse weight matrix $\boldsymbol{W}^{(l)}$ (i.e., $\boldsymbol{W}^{(l)}$ has only one nonzero entry) can also lead to an axis-aligned splitting rule.
Up to now, we have established the explicit connection between the ReLU network and previous axis-aligned (Gordon and Olshen, 1985; Wang et al., 2016) and non-axis-aligned (Jaeger et al., 2019) survival trees. Since the derived ReLU network itself is a neural network, its connection to other deep survival models (Katzman et al., 2018; Kvamme et al., 2019; Lee et al., 2018; Ren et al., 2019) is trivial.
Figure 2. The change of rank ratio of matrix $\boldsymbol{O}$ over epochs.
@@ 2.2. Statistically-Driven Topology Optimization
Although we have constructed a ReLU network whose behavior inherently resembles a tree, it is not specifically optimized for survival analysis. One notable limitation of the previously derived ReLU network in Sec. 2.1 is that its network topology is defined a priori. This implies that the tree constructed by the ReLU network is fully expanded (i.e., a fully-expanded dense tree).
As the activation patterns $\boldsymbol{O}$ uniquely encoded the partitioning of input space in the derived ReLU network, the topology of the learned tree structure can be reflected by $\boldsymbol{O}$. We observe that the activation patterns matrix $\boldsymbol{O}$ is typically low-ranked (see Fig. 2). This suggests that the learned ReLU network, however, produces a sparse tree topology: some partitions/nodes should be pruned. Based on this observation, we employ the log-rank test to localize which node should be pruned. Specifically, at each forward pass, we can reconstruct the tree structure of the derived ReLU network recursively (see Fig. 1(c)). Simultaneously, we can perform the log-rank test at each splitting node and merge the nodes whose p-value from the log-rank test is less than 0.05 ($p<0.05$) by setting the corresponding activation patterns to 0. This enables us to dynamically optimize the derived ReLU network’s topology and automatically prune the resulting tree structures.
However, when to perform the proposed topology optimization remains an open challenge. Model performance may deteriorate when optimizing network topology after training (Molchanov et al., 2016). However, optimizing network topology at early training iterations may restrict its representational capacity (Blalock et al., 2020). Again, we can leverage the rank of the activation patterns $\boldsymbol{O}$ to decide when to optimize the network topology. To achieve this, we keep track of the changes in the rank of the matrix $\boldsymbol{O}$ over time. Once the topology of the learned tree structures is unchanged, we optimize the network topology.
@@ 2.3. Model Learning
Complexity analysis. Compared to standard MLP, SurvReLU computes the derivative of each neuron w.r.t. $\boldsymbol{x}$, which requires an additional back-propagation ($\mathcal{O}(L^{3})$). By leveraging the dynamic programming trick as outlined in (Lee and Jaakkola, 2020), this can be reduced to $\mathcal{O}(L^{2})$. The log-rank test also introduces additional computation, which will take $\mathcal{O}(N\log N)$ to sort the time $T$. However, this can be handled once at the very beginning of each training iteration.
@@
4:     reconstruct the tree node at the $l$-th layer (see Fig. 1(c))
5:     log-rank test on two branches ($\boldsymbol{z^{(l)}}\geq 0$ and $\boldsymbol{z^{(l)}}<0$)
@@
{NiceTabular}
l—c—c—c—c[colortbl-like] Methods Simulated Linear Simulated Gaussian SUPPORT METABRIC
@@
3. Experiments
@@
We validated SurvReLU on two simulated and two real datasets. Datasets. The simulated datasets were sourced from (Katzman et al., 2018), consisting of a training ($N=4000$), validation ($N=1000$), and testing ($N=1000$) set. Each sample includes $d=10$ covariates (i.e., $x_{0},x_{1},\cdots,x_{9}$) drawn from a uniform distribution on $[-1,1)$. The event time $T$ was generated based on an exponential Cox model. Further information about the simulation process can be found in (Katzman et al., 2018). These simulated datasets encompass two subcategories characterized by either a linear risk function (i.e., $x_{1}$: $h(x)=x_{0}+2x_{1}$) or a Gaussian risk function (i.e., $h(\boldsymbol{x})=\log(r_{max})\exp(-(x_{0}^{2}+x_{1}^{2})/2\sigma^{2})$; $r_{max}=5.0$ and $\sigma=0.5$). Additionally, two benchmark real datasets with censorship of $32\%$ and $42\%$ were used: (i) SUPPORT dataset, originating from the Study to Understand Prognoses Preference Outcomes and Risks of Treatment (Knaus et al., 1995). It comprises 8873 patients (patients with missing features were dropped) with 14 covariates. (ii) METABRIC dataset is designed to improve breast cancer treatment recommendations (Curtis et al., 2012), including gene expression and clinical data from 1,980 patients. The data was processed to integrate four gene indicators and five clinical features, resulting in 9 covariates.
@@
Evaluation metrics. The evaluation of the performance for survival analysis was carried out by the time-dependent concordance index ($C^{td}_{\tau}$). Specifically, C-index is designed to assess a model’s ability to predict the ordering of event times through ranking.
@@
Experimental setup. (i) On the simulated datasets, we adhered to the default training, validation, and testing data splits outlined in (Katzman et al., 2018). All the covariates were z-scored to have a zero mean and unit variance. The 95% confidence intervals were obtained by bootstrapping (Katzman et al., 2018). (ii) We conducted five-fold cross-validation for the SUPPORT and METABRIC datasets to thoroughly assess the performance and generalizability of all methods. The mean ($\pm$ standard deviation) of C-index is reported. For preprocessing, we applied one-hot encoding to all categorical variables and z-score for all numerical variables. (iii) The hyperparameter tuning for the proposed model (i.e., learning rate, batch size, and $L$) was carried out by the random search (Bergstra and Bengio, 2012). The optimal learning rate was found to be $0.1$. The batch size was set to 1024. We recommend a large batch size to ensure stable statistics in the log-rank test. The optimal number of layers was [6, 30, 16, 14] for simulated Linear, simulated Gaussian, SUPPORT, and METABRIC, respectively. All the experiments were implemented in PyTorch and performed on a Nvidia GTX 1080Ti.
4. Results
@@
Main results. SurvReLU competed favorably with other deep survival and tree-based survival models on both simulated and real datasets in terms of $C^{td}$ (Table 2.3). We observed that (i) on the simulated datasets, SurvReLU (Cont.) and SurvReLU (Disc.) achieved similar performance but still outperformed other competing methods. Remarkably, SurvReLU approximated true Linear/Gaussian risk functions as effectively as other deep survival models (see Fig. 3(a)). However, the axis-aligned and non-axis-aligned tree-based survival models showed inferior performance in modeling the risk function. This suggests that SurvReLU retains the representational capacity of a neural network. Although SurvReLU has a non-axis-aligned partitioning rule similar to ORSF, we conjecture that the superiority of SurvReLU over other tree-based survival models arises from its end-to-end optimization instead of greedy expansion (e.g., fitting a CPH at each split in ORSF). (ii) On the real datasets, SurvReLU (Cont.) outperformed other competing continuous-time deep survival models (DeepSurv and Cox-CC). SurvReLU (Disc.) outperformed discrete-time DeepHit and other tree-based survival models, even without model ensembles. Importantly, SurvReLU resulted in a compact tree structure that was inherently as interpretable as tree-based survival models (Fig. 3(b) and (c)), a feature none of the other deep survival models can provide.
@@
Ablation studies. We conducted ablation on the number of layers $L$ in a SurvReLU network as well as the sparsity of the weight matrices $\boldsymbol{W}=\{\boldsymbol{W}^{(1)},\cdots,\boldsymbol{W}^{(L)}\}$. The sparsity of $\boldsymbol{W}$ is enforced by recursively soft-thresholding $\boldsymbol{W}$ with a strength $\lambda$ (see e.g. (Donoho, 1995; Kusupati et al., 2020; Xiao et al., 2023; Qiu et al., 2023)). We observed that (i) similar to many deep survival models, the performance of SurvReLU typically saturated after reaching a certain number of layers (e.g., $L=20$ in Fig. 4(Left)). (ii) The performance of SurvReLU was traded-off with the sparsity of $\boldsymbol{W}$ and hence the interpretability of the model (Fig. 4(Right)). This is because a more sparse $\boldsymbol{W}$ pushes SurvReLU to result in an axis-aligned tree structure (see Sec. 2.1), which is better to interpret.
@@
Figure 4. Ablations on number of layers $L$ (Left) and sparsity of $\boldsymbol{W}$ (Right) using the simulated Gaussian dataset.